# Is the Interstellar Object 3I/ATLAS Alien Technology?

### v1

@@ 1 Introduction
The hypothesis in question is that the recent interstellar visitor to our Solar System, 3I/ATLAS (Seligman et al., 2025; Bolin et al., 2025; Opitom et al., 2025; Alvarez-Candal et al., 2025; Hopkins et al., 2025, 2024; Taylor & Seligman, 2025; Kakharov & Loeb, 2025; Loeb, 2025a, b), is a technological artifact, and furthermore has active intelligence. If this is the case, then two possibilities follow: first that its intentions are entirely benign and second they are malign.
In the first case, humanity need do nothing save await the arrival of this intelligence with open arms. It is the second eventuality which is of most concern, and according to the so-called ’Dark Forest’ resolution to the ’Fermi Paradox’, would be more likely, as it would neatly explain the singular lack of success of the SETI initiative to-date (SETI Institute, 2025).
Discovered on 1st July 2025, by the Asteroid Terrestrial-impact Last Alert System, 3I/ATLAS is, as its designation indicates, the latest interstellar object (or interloper) to be discovered passing through our Solar System. The first, 1I/’Oumuamua, detected in 2017, was only visible for a period of 2 months, though various anomalous features of this object have yet to be clarified (Bannister et al., 2019; Loeb, 2022). Despite this, there are still very entrenched opinions on the subject in the scientific community as Eldadi, Tenenbaum and Loeb submitted to ”Psychological Review” on July 9th, 2025 (Eldadi, O. and Tenenbaum, G. and Loeb, A., 2025).
@@ 3 Possible strategies and motivations
Mars -5.39E-05 -5.33E-04 5.14E-04
Mars 1.19E-04 -4.76E-04 3.55E-04
Jupiter 3.53E-05 2.64E-05 -3.85E-05

### v2

@@ 1 Introduction
The hypothesis in question is that the recent interstellar visitor to our Solar System, 3I/ATLAS (Seligman et al., 2025; Bolin et al., 2025; Opitom et al., 2025; Alvarez-Candal et al., 2025; Hopkins et al., 2025, 2024; Taylor & Seligman, 2025; Kakharov & Loeb, 2025; Loeb, 2025a, b), is a technological artifact, and furthermore has active intelligence. If this is the case, then two possibilities follow: first that its intentions are entirely benign and second they are malign, or somewhere inbetween.
To address the extreme cases in turn, in the first case, humanity need do nothing save await the arrival of this intelligence with open arms. It is the second eventuality which is of most concern, and according to the so-called ’Dark Forest’ resolution to the ’Fermi Paradox’, would be more likely, as it would neatly explain the singular lack of success of the SETI initiative to-date (SETI Institute, 2025).
Discovered on 1st July 2025, by the Asteroid Terrestrial-impact Last Alert System, 3I/ATLAS is, as its designation indicates, the latest interstellar object (or interloper) to be discovered passing through our Solar System. The first, 1I/’Oumuamua, detected in 2017, was only visible for a period of 2 months, though various anomalous features of this object have yet to be clarified (Bannister et al., 2019; Loeb, 2022). Despite this, there are still very entrenched opinions on the subject in the scientific community as Eladi, Tenenbaum and Loeb submitted to ”Psychological Review” on July 9th, 2025 (O. Eldadi and G. Tenenbaum and A. Loeb, 2025).
@@ 3 Possible strategies and motivations
Mars -5.39×$\times 10^{-5}$5.33×$\times 10^{-4}$ 5.14×$\times 10^{-4}$
Mars 1.19×$\times 10^{-4}$4.76×$\times 10^{-4}$ 3.55×$\times 10^{-4}$
Jupiter 3.53×$\times 10^{-5}$ 2.64×$\times 10^{-5}$3.85×$\times 10^{-5}$
@@ 4 Non-gravitational Accelerations
5 Jupiter Oberth and Aero-Capture
Jupiter Label Units Value
Gravitational Parameter GM $\mathrm{k}\mathrm{m}^{3}\,\mathrm{s}^{-2}$ 1.27×$\times 10^{8}$
Mean Equatorial Radius Re $\mathrm{k}\mathrm{m}$ 71492
Mean Polar Radius Rp $\mathrm{k}\mathrm{m}$ 66854
J2 Harmonic J2 N/A 0.014736
Period s 35730
Table 4: Pertinent Jupiter parameters from NTRS - NASA Technical Report Service (2021)
While 3I/ATLAS might be a circular sail, another option, with the advantage of a high degree of symmetry is a sphere. To maintain rigidity and strength as well as a very low areal mass density led previous researchers to propose interstellar spherical sails made of either graphene (Heller et al., 2017; Heller & Hippke, 2017; Matloff, 2013), or Aerographite (Heller et al., 2020; Karlapp et al., 2024). For this discussion we shall assume multi-layer graphene, with an areal mass density of 0.758×$\times 10^{-7}$ $\mathrm{k}\mathrm{g}\,\mathrm{m}^{-2}$, and an absorptivity of 0.02. To achieve 3I/ATLAS’s observed albedo of 0.04 (i.e. an absorptivity of 0.96) means at least 96 layers. With a radius of 10 $\mathrm{k}\mathrm{m}$, 3I/ATLAS’s outer layer masses at least 91.4 tonnes. Assuming the normalised acceleration advised above, 3I/ATLAS acting as an absorbing Sail, masses at most 1,160 tonnes. Thus it can contain multiple internal layers to brace it against outside forces as well as significant payload, like sub-probes and sensor equipment. The implied cross-sectional mass-density is 0.00369 $\mathrm{k}\mathrm{g}\,\mathrm{m}^{-2}$.
At Jupiter, 3I/ATLAS has several options one of which is Aero-Capture into a Highly Elliptical Jupiter Orbit (HEJO). Using OITS to model the orbit gives a Hyperbolic Excess of 65.5913 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$ relative to Jupiter’s centre. Assuming the current physical parameters for Jupiter (NTRS - NASA Technical Report Service, 2021) (See Table 4) 3I/ATLAS will be moving at 88.33 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$ when encountering the atmosphere at an altitude of $\sim{899}$ $\mathrm{k}\mathrm{m}$ above the 1 bar reference level. Jupiter sidereal rotation period of 35,730 seconds gives an Equatorial rotational speed of 12.729 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$ at that altitude. Therefore the Entry Speed is 75.6 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$ relative to the atmosphere.
For a capture orbit with an Apo-Jove of 527 Jupiter radii, the Peri-Jove speed needs to be 58.74 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$, or 46 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$ relative to the atmosphere. These are just first pass estimates, to give some idea of the re-entry conditions. Assuming a constant braking in a path equal to one Jupiter radius in length (71.5×$\times 10^{6}$ $\mathrm{m}$), the deceleration for Aero-Capture is just  25 $\mathrm{m}\,\mathrm{s}^{-2}$. As 3I/ATLAS is above circular orbital velocity for the whole event, the centrifugal acceleration it experiences is counter to gravity, unlike a landing re-entry which must maintain positive lift to maximize braking distance at high altitude.
With such a low areal mass-density of 3.69×$\times 10^{-3}$ $\mathrm{k}\mathrm{g}\,\mathrm{m}^{-2}$, the energy dissipation is 7.01 $\mathrm{k}\mathrm{W}\,\mathrm{m}^{-2}$ – high, but not the multi-megawatt levels experienced during an Apollo mission Lunar re-entry to Earth (NTRS - NASA Technical Report Service, 2013). Additionally the total heat load is 6.64 $\mathrm{M}\mathrm{J}\,\mathrm{m}^{-2}$, much lower than the 426.5 $\mathrm{M}\mathrm{J}\,\mathrm{m}^{-2}$ experienced by Apollo. Once decelerated into orbit, conceivably 3I/ATLAS can interact with Jupiter’s Galilean moons and magnetic field to maneuver without use of propellant.
Covert reconnaissance of an inhabited star system is described in some detail in Stanislaw Lem’s influential fictional critique of SETI, “Fiasco” (Lem, 1987). In the novel Earth vehicle “Hermes” approaches an inhabited star system and adopts the guise of a parabolic comet, applying an artificial crust and ejecting gases. Later it uses close flybys of gas giants to decelerate and obscure its propulsive manoeuvering. “Hermes” adopts this cautiously covert entry into a star system due to observing possible hostile activities remotely spread across the target system, centred on the inhabited planet.